# Distributed Algorithms for Potential Problems

### v3

@@ 3.1 Preliminary definitions and results
A key notion in our algorithm is that of an improving set. As already described in the introduction, the main insight of our approach is that if a node $u$ is part of some sequential flipping process that obtains a globally correct solution, we can detect this in $u$’s $O(\operatorname{log}n)$ neighborhood. What we will find in $u$’s neighborhood is an improving set with a good improving ratio. We will later prove this formally, so to state our algorithm, we give the necessary definitions here.
@@
In this paragraph, we define and describe the MPX clustering algorithm, which is a key component of our algorithm, following the paper that first introduced it [MPX13].
@@
Consider any node $v\in V$, and take the node $u$ that realizes $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. By contradiction, suppose there exists a node $w$ such that $\operatorname{dist}_{\text{shift}}(u,u)>\operatorname{dist}_{\text{shift}}(w,u)$. Then, we have that $\lambda_{u}>\operatorname{dist}(w,u)-\lambda_{w}$. This implies that $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}>\operatorname{dist}(u,v)+\operatorname{dist}(u,w)-\lambda_{w}\geq\operatorname{dist}(w,v)-\lambda_{w}=\operatorname{dist}_{\text{shift}}(w,v)$ by the triangular inequality. Hence, $u$ does not realize the minimum of $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$, reaching a contradiction.
Now, we say that each node $v$ joins the cluster of the node $u$ that minimizes the shifting distance from $u$ to $v$, that is, $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. Let $d=O(\operatorname{log}n/\rho)$. In [MPX13], it is shown that this procedure produces a $(\rho,d)$-decomposition of $G$ with high probability. While the original procedure is defined for the PRAM model, [EN22] showed that it can be implemented in the LOCAL model through an $O(\operatorname{log}n/\rho)$-round randomized LOCAL algorithm.
Now let two nodes $v,v^{\prime}$, such that $\operatorname{dist}(v,v^{\prime})\leq k$. Let $u_{v}$ be the center of the cluster that $v$ joins, that is the node realizing $\min_{u\in V}\operatorname{dist}_{\text{shift}}(u,v)$. Let $d_{1}\leq\ldots\leq d_{n}$ be the distances from $v_{1}$ to any node $v_{1}=u_{1},\ldots,u_{n}\in V$. Consider $\lambda_{i}$ to be the exponential random variable associated to the node $u_{i}$. By Lemma 3.5, we have that, with probability $O(\rho k/2)$, it holds that the smallest value and the second-smallest value in the family $\{(d_{i}-\lambda_{i})\}_{i\in[n]}$ are within distance $k$ of each other. Let $k=1/(c\rho)$ for a large enough constant $c$ so that the probability $O(\rho k/2)$ is strictly less than $1/2$.
@@ 3.2 Algorithm description
Our algorithm simply repeats the following process $\operatorname{log}n$ times: Compute a clustering, such that for every node $v$ with constant probability its entire $O(\operatorname{log}^{4}n)$ neighborhood is contained within its cluster. In each cluster we brute force a maximal sequence of improving sets, each of improving ratio at least $R$. We then apply the relabeling implied by this sequence in every cluster and inform all nodes of the cluster. Informally, we now consider the nodes in these clusters to be ”safe”. To ensure that we do not ”undo” the progress we have made in previous iterations, we have to increase the minimum improving ratio $R$ between every phase by a small value (roughly $\varepsilon$).
@@
1:Number of nodes $n$, description of the LOP $(\Pi,\Psi)$.
@@
9:  Run MPX with $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)$ $\triangleright$ $d=O(\operatorname{log}n/\rho)$ by Lemma 3.6
@@
16:      2. The diameter of each $A_{i}$ is at most $c_{3}\operatorname{log}(n)/\varepsilon$
@@
For every labeled centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$, there is a way to reassign the labels to $v_{H}$ and its incident half-edges so that the potential $\operatorname{Pot}(G,\ell)$ decreases and the centered graph $(H,v_{H})$ with the new labeling belongs to $\mathcal{C}$. Let $\ell^{\prime}$ be the new labeling: $(\{v_{H}\},\ell^{\prime})$ is a minimal improving set with some positive improvement $\hat{\lambda}$. Let $\lambda$ be the minimum over all possible values of $\hat{\lambda}$. We say that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$ (which is initialized in Line 3). Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$.
@@
The idea of the algorithm is as follows. Set $\varepsilon=\lambda/(100c_{1}\operatorname{log}(n))$, where $c_{1}$ is a small enough constant (Line 5), and set a threshold improving ratio to be $R=\lambda/4$ (Line 4). We start with an arbitrary labeling of the nodes and half-edges of $G$. Then, we have $\Theta(\operatorname{log}n)$ phases (from Line 8 to Line 26), where in each phase we run the MPX clustering algorithm (Line 9) with parameters $\rho=10\cdot c_{2}\cdot\varepsilon^{2}/(c\operatorname{log}^{2}n)=O(\operatorname{log}^{4}n)$ and $d=O(\operatorname{log}n/\rho)$, where $c,c_{2}$ are large enough constants that come from Lemmas 3.6 and 4.8, respectively: the bound on $d$ then holds by Lemma 3.6. Each run of MPX takes $O(d)=O(\operatorname{log}^{5}n)$ rounds. We assume that each cluster $C$ given by the MPX algorithm comes with a leader node $v_{C}$, that can be the center of the cluster itself. This anyway can be obtained in $O(d)$ rounds after each MPX run. Each leader node looks at the subgraph $H$ of $G$ induced by the cluster $C$ it belongs to, and it computes a maximal sequence of $R$-improving sets $(A_{i},\ell_{i})_{i=1}^{k}$ in $H$ w.r.t. $\ell$. The sequence satisfies that $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$ (from Line 13 to Line 14), ensuring that the sequences of different clusters do not interfere with each other. Furthermore, we require the diameter of each set in the sequence to be at most $c_{3}\operatorname{log}n/\varepsilon$ (the choice of the large constant $c_{3}$ comes from Lemma 4.5). When we say that the sequence is maximal, we mean that in $H$ labeled by $\ell_{k}$ there is no minimal improving set $(A^{\prime},\ell^{\prime})$ such that $\operatorname{IR}(A^{\prime},\ell_{k},\ell^{\prime})>R$ and $\mathcal{N}_{2r+1}[A^{\prime}]\subseteq H$. Note that such a sequence must exist and terminate, as the potential of $H$ is always reduced after every relabeling (and cannot be reduced arbitrarily): furthermore, the sequence can be computed in $O(d)$ rounds by brute-force. For this the leader simply collects the topology and labeling of its entire cluster and brute forces every possible sequence in local computation. Then, $v_{C}$ adopts the labeling defined by $\ell_{k}$ and also broadcasts $\ell_{k}$ to every node of the cluster $C$ (Line 17). Every non-leader node just waits for the leader node to broadcast the new labeling (lines 20 and 21), and then it adopts the new labeling (Line 22). Then, the phase terminates and the nodes update the reference value of the improving ratio $R$ to $R+\lambda/(20c_{1}\operatorname{log}n)$ (Line 25). The algorithm terminates after $\Theta(\operatorname{log}n)$ phases, and it returns the labels of the nodes and of the incident half-edges (Line 27).
@@
$T(n)=O(\operatorname{log}n\cdot d)=O(\operatorname{log}^{4}n/\varepsilon^{2})=O(\operatorname{log}^{6}n).$
In the rest of the paper, we prove the following theorem.
@@
Let $(\Pi,\Psi)$ be a LOP as defined in Definition 2.7. The randomized LOCAL algorithm described in Algorithm 1 solves $(\Pi,\Psi)$ in $O(\operatorname{log}^{6}n)$ rounds w.h.p.
@@ 4 Analysis of the algorithm
For the rest of the section, we fix an LOP of interest $(\Pi,\Psi)$ and any input graph $G=(V,E)$ of $n$ nodes that is already $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled (hence it has an input and a – possibly invalid – output labeling). Let $\ell$ denote the labeling of $G$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$, while $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$. Trivially, $\lambda\leq\Lambda$. We remark that all the values $r,\Delta,\lambda,\Lambda$ are constant values and will be hidden in the asymptotic notation. In the proofs of the lemmas in Section 4.1, we do keep track of a loose dependency in these parameters, which is polynomial. The reader might extract a $\text{poly}(\Delta^{r},\Lambda)\cdot\text{polylog}(n)$-time algorithm from this analysis; however, we remark that we did not optimize such dependencies.
@@
$\ell^{\star}_{A}=\arg\max_{x\in\mathscr{L}}\set{\operatorname{Imp}(A,\ell,x)}$,
@@
The whole analysis of the algorithm relies on the following two lemmas (Lemmas 4.2 and 4.4), the last of which requires involved technical claims. The first lemma states that each phases terminates.
@@
At the end of phase $i$ of the algorithm, all minimal improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ and improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are fully contained in the set
@@
for $t_{1}=c_{3}\operatorname{log}n/\varepsilon+10(r+1)$ and $t_{2}=c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+t_{1}$, where $c_{2}$ and $c_{3}$ are as in Lines 9 and 16, respectively.
The proof of Lemma 4.4 relies on the following two lemmas. The first one states that for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. The proof of Lemma 4.5 is involved, and we defer it to Section 4.1.
@@
Let $\ell$ be a labeling of $G=(V,E)$ and let $(S,\ell_{S})$ be a $\beta$-minimal improving set, for some $\beta>0$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for every $0<\varepsilon<\min{\{\beta,\lambda\}}$ and any $v\in S$ there exists a large enough constant $c_{3}$, a $k\leq c_{3}\operatorname{log}n/\varepsilon$, and a subset $A\subseteq S\cap\mathcal{N}_{k}[v]$, such that $\operatorname{IR}(A,\ell,\ell_{A}^{\star})\geq\beta-\varepsilon$. Furthermore, $(A,\ell_{A}^{\star})$ is minimal.
@@
Let $(A_{i},\ell_{i})_{i=1}^{h}$ be a sequence of $\beta$-improving sets of diameter at most $c_{3}\operatorname{log}n/\varepsilon$, for some $\beta>0$ and $0<\varepsilon<\min{\{\beta,\lambda\}}$, where $c_{3}$ is the constant given by Lemma 4.5, and $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for each $i\in[h]$ and each $v\in A_{i}$, there exists a large enough constant $c_{2}$ and an improving set $(A,\ell_{A})$ of diameter $c_{3}\operatorname{log}n/\varepsilon$, such that $\operatorname{IR}(A,\ell,\ell_{A})\geq\beta-2\varepsilon$, that is fully contained in the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap(\cup_{j\in[h]}A_{j})$. Furthermore, $(A,\ell_{A})$ is minimal.
@@
For any phase $i>0$, let $\mathcal{C}^{(i)}_{1},\ldots,\mathcal{C}^{(i)}_{h_{i}}$ be the clusters determined by MPX at phase $i$ (Line 9), let $\ell_{i}$ be the labeling of $G$ at the end of phase $i$ (Lines-18 and 23), and let $R_{i}$ be the improving ratio initialized at the end of phase i1i-1 (Line 25), with $R_{1}=R$ as in Line 4. We proceed by induction on the phase $i$.
If $i=1$, $R_{1}=R$ as initialized at Line 4. A minimal improving set $(A,\ell_{A})$ of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ with $\operatorname{IR}(A,\ell_{1},\ell_{A})\geq R_{1}$ must be such that there is no $s\in[h_{1}]$ such that $\mathcal{N}_{2r+1}[A]\subseteq C^{(1)}_{s}$, otherwise we are breaking maximality in Line 14. Hence, $\operatorname{dist}(A,B_{1})\leq 2r$ and we have that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{1}]$, as the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$. Now, assume $i>1$ and the statement to be true for all phases $j=1,\ldots,i-1$. Let $(A,\ell_{A})$ be a minimal improving set of diameter at most $c_{3}\operatorname{log}n/\varepsilon$ such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}$. For the same reason as in the case $i=1$, it holds that $A\subseteq\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon+2r}[B_{i}]$. Suppose, by contradiction, that $A$ is not contained in

### v4

@@ 3.1 Preliminary definitions and results
A key notion in our algorithm is that of an improving set. As already described in the introduction, the main insight of our approach is that if a node $u$ is part of some sequential flipping process that obtains a globally correct solution, we can detect this in $u$’s $O(\operatorname{log}n)$ neighborhood (this only holds for $\Delta=O(1)$, but we will consider the more general case). What we will find in $u$’s neighborhood is an improving set with a good improving ratio. We will later prove this formally, so to state our algorithm, we give the necessary definitions here.
@@
In this paragraph, we define and describe the MPX clustering algorithm, which is a key component of our algorithm, following the paper that first introduced it [40].
@@
Consider any node $v\in V$, and take the node $u$ that realizes $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. By contradiction, suppose there exists a node $w$ such that $\operatorname{dist}_{\text{shift}}(u,u)>\operatorname{dist}_{\text{shift}}(w,u)$. Then, we have that $\lambda_{u}>\operatorname{dist}(w,u)-\lambda_{w}$. This implies that $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}>\operatorname{dist}(u,v)+\operatorname{dist}(u,w)-\lambda_{w}\geq\operatorname{dist}(w,v)-\lambda_{w}=\operatorname{dist}_{\text{shift}}(w,v)$ by the triangle inequality. Hence, $u$ does not realize the minimum of $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$, reaching a contradiction.
Now, we say that each node $v$ joins the cluster of the node $u$ that minimizes the shifting distance from $u$ to $v$, that is, $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. Let $d=O(\operatorname{log}n/\rho)$. In [40], it is shown that this procedure produces a $(\rho,d)$-decomposition of $G$ with high probability. While the original procedure is defined for the PRAM model, [28] showed that it can be implemented in the LOCAL model through an $O(\operatorname{log}n/\rho)$-round randomized LOCAL algorithm.
Now let two nodes $v,v^{\prime}$, such that $\operatorname{dist}(v,v^{\prime})\leq k$. Let $u_{v}$ be the center of the cluster that $v$ joins, that is the node realizing $\min_{u\in V}\operatorname{dist}_{\text{shift}}(u,v)$. Let $d_{1}\leq\ldots\leq d_{n}$ be the distances from $v$ to the nodes $u_{1},\ldots,u_{n}\in V$. Consider $\lambda_{i}$ to be the exponential random variable associated to the node $u_{i}$. By Lemma 3.5, we have that, with probability $O(\rho k/2)$, it holds that the smallest value and the second-smallest value in the family $\{(d_{i}-\lambda_{i})\}_{i\in[n]}$ are within distance $k$ of each other. Let $k=1/(c\rho)$ for a large enough constant $c$ so that the probability $O(\rho k/2)$ is strictly less than $1/2$.
@@ 3.2 Algorithm description
Our algorithm simply repeats the following process $\operatorname{log}n$ times: Compute a clustering, such that for every node $v$ with constant probability its entire radius-$O(\Delta^{2r}\Lambda^{2}\operatorname{log}^{4}(\Lambda n/\lambda))$ neighborhood is contained within its cluster, where $\Lambda$ and $\lambda$ are later defined in this section, and are related to the description of the specific LOP of interest. In each cluster we brute-force a maximal sequence of improving sets, each of improving ratio at least $R$. We then apply the relabeling implied by this sequence in every cluster and inform all nodes of the cluster. Informally, we now consider the nodes in these clusters to be ”safe”. To ensure that we do not ”undo” the progress we have made in previous iterations, we have to increase the minimum improving ratio $R$ between every phase by a small value (roughly $\varepsilon$).
@@
1:Number of nodes $n$, max-degree $\Delta$, description of the LOP $(\Pi,\Psi)$.
@@
9:  Run MPX with $\rho=\varepsilon^{2}/(10c\cdot c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda))$ $\triangleright$ $d=O(\operatorname{log}n/\rho)$ by Lemma 3.6
@@
16:      2. The diameter of each $A_{i}$ is at most $c_{3}\Delta^{r}\Lambda\operatorname{log}(n)/\varepsilon$
@@
For every labeled centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$ that does not belong to $\mathcal{C}$, there is a way to reassign the labels to $v_{H}$ and its incident half-edges so that the potential $\operatorname{Pot}(G,\ell)$ decreases and the centered graph $(H,v_{H})$ with the new labeling belongs to $\mathcal{C}$. Let $\ell^{\prime}$ be the new labeling: $(\{v_{H}\},\ell^{\prime})$ is a minimal improving set with some positive improvement $\hat{\lambda}$. Let $\lambda$ be the minimum over all possible values of $\hat{\lambda}$. We say that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$ (which is initialized in Line 3). Similarly, let $\Lambda$ be the maximum possible decrease of the potential over any centered graph $(H,v_{H})\in\mathcal{L}(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}},2r,\Delta)$. We say that $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$. Note that, given the description of the LOP and the value of $\Delta$, the nodes of the graph can compute the values $\lambda$ and $\Lambda$ locally.
@@
The idea of the algorithm is as follows. Set $\varepsilon=\lambda/(100c_{1}\operatorname{log}(n))$, where $c_{1}$ is a large enough constant (Line 5), and set a threshold improving ratio to be $R=\lambda/4$ (Line 4). We start with an arbitrary labeling of the nodes and half-edges of $G$. Then, we have $\Theta(\operatorname{log}n)$ phases (from Line 8 to Line 26), where in each phase we run the MPX clustering algorithm (Line 9) with parameters $\rho=\varepsilon^{2}/(10c\cdot c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda))=O(\lambda^{2}/(\Delta^{2r}\Lambda^{2}\operatorname{log}^{4}(\Lambda n/\lambda)))$ and $d=O(\operatorname{log}n/\rho)$, where $c,c_{2}$ are large enough constants that come from Lemmas 3.6 and 4.8, respectively: the bound on $d$ then holds by Lemma 3.6. Each run of MPX takes $O(d)=O(\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{5}(\Lambda n/\lambda))$ rounds. We assume that each cluster $C$ given by the MPX algorithm comes with a leader node $v_{C}$, that can be the center of the cluster itself. This anyway can be obtained in $O(d)$ rounds after each MPX run. Each leader node looks at the subgraph $H$ of $G$ induced by the cluster $C$ it belongs to, and it computes a maximal sequence of $R$-improving sets $(A_{i},\ell_{i})_{i=1}^{k}$ in $H$ w.r.t. $\ell$. The sequence satisfies that $\mathcal{N}_{2r+1}[A_{i}]\subseteq C$ for all $i\in[k]$ (from Line 13 to Line 14), ensuring that the sequences of different clusters do not interfere with each other. Furthermore, we require the diameter of each set in the sequence to be at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ (the choice of the large constant $c_{3}$ comes from Lemma 4.5). When we say that the sequence is maximal, we mean that in $H$ labeled by $\ell_{k}$ there is no minimal improving set $(A^{\prime},\ell^{\prime})$ such that $\operatorname{IR}(A^{\prime},\ell_{k},\ell^{\prime})>R$ and $\mathcal{N}_{2r+1}[A^{\prime}]\subseteq H$. Note that such a sequence must exist and terminate, as the potential of $H$ is always reduced after every relabeling (and cannot be reduced arbitrarily): furthermore, the sequence can be computed in $O(d)$ rounds by brute-force. For this the leader simply collects the topology and labeling of its entire cluster and brute forces every possible sequence in local computation. Then, $v_{C}$ adopts the labeling defined by $\ell_{k}$ and also broadcasts $\ell_{k}$ to every node of the cluster $C$ (Line 17). Every non-leader node just waits for the leader node to broadcast the new labeling (Lines 20 and 21), and then it adopts the new labeling (Line 22). Then, the phase terminates and the nodes update the reference value of the improving ratio $R$ to $R+\lambda/(20c_{1}\operatorname{log}n)$ (Line 25). The algorithm terminates after $\Theta(\operatorname{log}n)$ phases, and it returns the labels of the nodes and of the incident half-edges (Line 27).
@@
$T(n)=O(\operatorname{log}n\cdot d)=O(\operatorname{log}n\cdot\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{5}(\Lambda n/\lambda))=O(\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda)).$
We remind the reader that $r$ is a constant, while the values $\lambda,\Lambda$ can depend on the maximum degree $\Delta$ of the input graph: if $\Delta=O(1)$, then $\Lambda=O(1)$ and $\lambda=O(1)$, disappearing from asymptotic notation. In the rest of the paper, we prove the following theorem.
@@
Let $(\Pi,\Psi)$ be a LOP as defined in Definition 2.7. The randomized LOCAL algorithm described in Algorithm 1 solves $(\Pi,\Psi)$ in $O(\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda))$ rounds w.h.p. If $\Delta=O(1)$, then the running time of Algorithm 1 is $O(\operatorname{log}^{6}(n))$ rounds.
Note that, for locally optimal cut, it holds that $r=1$, $\lambda=1$, and $\Lambda=\Delta$. Since $\Delta\leq n$, by Theorem 3.7, we would obtain a randomized runtime of $O(\Delta^{4}\operatorname{log}^{6}(n))$ rounds. However, as we argue in the appendix, for a natural subclass of LOPs, called Locally Optimal Problems on Edges (LOPEs), we can do better and obtain a better runtime (see Definitions A.2 and A.11). Note that the proof scheme of Theorem A.11 is essentially the same as the one of Theorem 3.7 and follows the same lines: it just exploits some additional properties of LOPEs to get better inequalities. The related result for locally optimal cut is the following.
Corollary 3.8.
The randomized LOCAL algorithm described in Algorithm 2 solves the locally optimal cut problem in $O(\Delta^{2}\operatorname{log}^{6}(n))$ rounds w.h.p.
Proof.
Just apply Theorem A.11 to the locally optimal cut problem, which is a special case of LOPE with $\lambda=1$. Note that there is no need to rescale the potential function (as done in Appendix A) of the locally optimal cut problem, as $\Gamma=1$ already. ∎
3.3 Derandomization
The algorithm described in Algorithm 1 can be derandomized using standard techniques.
Let us define the notion of a $(\alpha,d)$-network decomposition of a graph $G=(V,E)$.
Definition 3.9 (Network decomposition).
A $(\alpha,d)$-network decomposition of a graph $G=(V,E)$ is a partition of the nodes $V$ into sets $C_{1},\ldots,C_{k}$, which we call clusters, such that:
  • For every $i$, the strong diameter of each $C_{i}$ is at most $d$.
  • The clusters can be colored with $\alpha$ colors so that no two adjacent clusters share the same color.
From [35, 33, 7, 32], the following holds.
Theorem 3.10.
Suppose we are given an $R(n)$-rounds randomized LOCAL algorithm $\mathcal{A}$ that solves a labeling problem $\Pi$ with high probability. Then, there exists a deterministic LOCAL algorithm that solves $\Pi$ that runs in time
$D(n)=O\left(R(n)\mathrm{ND}(n)+R(n)\alpha(n)d(n)\right),$
where $\mathrm{ND}(n)$ is the deterministic complexity of finding an $(\alpha(n),d(n))$-network decomposition of the input graph.
[32] proved the following theorem.
Theorem 3.11 ([32]).
There exists an $\tilde{O}(\operatorname{log}^{2}n)$-round deterministic LOCAL algorithm that computes a $(O(\operatorname{log}n),O(\operatorname{log}n))$-network decomposition of a graph $G=(V,E)$ of $n$ nodes.
We immediately obtain the following corollary.
Corollary 3.12.
There exists a deterministic LOCAL algorithm that solves any LOP $(\Pi,\Psi)$ in $O(\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda)\operatorname{log}^{2}(n)\operatorname{poly}(\operatorname{log}\operatorname{log}n))$ rounds. If $\Delta=O(1)$, then the running time of the deterministic algorithm is $\tilde{O}(\operatorname{log}^{8}(n))$ rounds.
As for the specific case of locally optimal cut, by applying the derandomization technique to the algorithm from Corollary 3.8, we have the following.
Corollary 3.13.
There exists a deterministic LOCAL algorithm that solves the locally optimal cut problem in $O(\Delta^{2}\operatorname{log}^{8}(n)\operatorname{poly}(\operatorname{log}\operatorname{log}n))$ rounds. If $\Delta=O(1)$, then the running time of the deterministic algorithm is $\tilde{O}(\operatorname{log}^{8}(n))$ rounds.
@@ 4 Analysis of the algorithm
For the rest of the section, we fix an LOP of interest $(\Pi,\Psi)$ and any input graph $G=(V,E)$ of $n$ nodes that is already $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled (hence it has an input and a – possibly invalid – output labeling). Let $\ell$ denote the labeling of $G$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$, while $\Lambda$ is the maximum improvement associated to $(\Pi,\Psi)$. Trivially, $\lambda\leq\Lambda$.
@@
$\ell^{\star}_{A}=\arg\max_{x\in\mathscr{L}}\{\operatorname{Imp}(A,\ell,x)\},$
@@
The whole analysis of the algorithm relies on the following two lemmas (Lemmas 4.2 and 4.4), the last of which requires involved technical claims. The first lemma states that each phase terminates.
@@
At the end of phase $i$ of the algorithm, all minimal improving sets of diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ and improving ratio at least $R_{i}$ (the value of $R$ initialized at the end of phase i1i-1 at Line 25) are fully contained in the set
@@
for $t_{1}=c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon+10(r+1)$ and $t_{2}=c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}+t_{1}$, where $c_{2}$ and $c_{3}$ are as in Lines 9 and 16, respectively.
The proof of Lemma 4.4 relies on the following two lemmas. The first one states that for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\Delta^{r}\Lambda\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. The proof of Lemma 4.5 is involved, and we defer it to Section 4.1.
@@
Let $\ell$ be a labeling of $G=(V,E)$ and let $(S,\ell_{S})$ be a $\beta$-minimal improving set, for some $\beta>0$. Recall that $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for every $0<\varepsilon<\min{\{\beta,\lambda\}}$ and any $v\in S$ there exists a large enough constant $c_{3}$, a $k\leq c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$, and a subset $A\subseteq S\cap\mathcal{N}_{k}[v]$, such that $\operatorname{IR}(A,\ell,\ell_{A}^{\star})\geq\beta-\varepsilon$. Furthermore, $(A,\ell_{A}^{\star})$ is minimal.
@@
Let $(A_{i},\ell_{i})_{i=1}^{h}$ be a sequence of $\beta$-improving sets of diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$, for some $\beta\geq\lambda$ and $0<\varepsilon<\lambda$, where $c_{3}$ is the constant given by Lemma 4.5, and $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Then, for each $i\in[h]$ and each $v\in A_{i}$, there exists a large enough constant $c_{2}$ and an improving set $(A,\ell_{A})$ of diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$, such that $\operatorname{IR}(A,\ell,\ell_{A})\geq\beta-2\varepsilon$, that is fully contained in the set $\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[v]\cap(\cup_{j\in[h]}A_{j})$. Furthermore, $(A,\ell_{A})$ is minimal.
@@
For any phase $i>0$, let $\mathcal{C}^{(i)}_{1},\ldots,\mathcal{C}^{(i)}_{h_{i}}$ be the clusters determined by MPX at phase $i$ (Line 9), let $\ell_{i}$ be the labeling of $G$ at the end of phase $i$ (Lines 18 and 23), and let $R_{i}$ be the improving ratio initialized at the end of phase i1i-1 (Line 25), with $R_{1}=R$ as in Line 4. We proceed by induction on the phase $i$.
If $i=1$, $R_{1}=R$ as initialized at Line 4. A minimal improving set $(A,\ell_{A})$ of diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ with $\operatorname{IR}(A,\ell_{1},\ell_{A})\geq R_{1}$ must be such that there is no $s\in[h_{1}]$ such that $\mathcal{N}_{2r+1}[A]\subseteq C^{(1)}_{s}$, otherwise we are breaking maximality in Line 14. Hence, $\operatorname{dist}(A,B_{1})\leq 2r$ and we have that $A\subseteq\mathcal{N}_{c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon+2r}[B_{1}]$, as the diameter of $A$ is at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$. Now, assume $i>1$ and the statement to be true for all phases $j=1,\ldots,i-1$. Let $(A,\ell_{A})$ be a minimal improving set of diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ such that $\operatorname{IR}(A,\ell_{i},\ell_{A})\geq R_{i}$. For the same reason as in the case $i=1$, it holds that $A\subseteq\mathcal{N}_{c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon+2r}[B_{i}]$. Suppose, by contradiction, that $A$ is not contained in